Circular No. 8140

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003en–2003ev Several additional apparent supernovae have been reported. Objects

2003en and 2003eq–2003eu were reported by L. G. Strolger, Space Telescope Science Institute (on behalf of the GOODS Treasury Team and the Hubble Higher-z Supernova Team; cf. IAUC 8125), from ‘Hubble Deep Field North’ ACS images, and each was detected in four z-band, two i-band, and two v-band images taken within 45 min of each other (all ACS magnitudes below are z-band). SNe 2003eo was reported by W. M. Wood-Vasey et al. (cf. IAUC 8130) from unfiltered NEAT images. SNe 2003ep and 2003ev were reported by M. Papenkova, D. Weisz, and W. Li from unfiltered LO-TOSS/KAIT images (cf. IAUC 8135). Each offset given below is that from the center of the nearest visible galaxy to the apparent supernova.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003en Jan. 3.2 12^{h}36^{m}33.^{s}12 +62^{o}13′48′′.1 25.5 0′′.07 W, 0′′.1 N
2003eo Mar. 27.14 11 03 09.19 − 5 46 09.4 20.2 1′′.5 W, 0′′.5 S
2003ep May 17.5 21 21 07.34 +23 05 05.8 15.7 3′′.9 W
2003eq May 24.7 12 37 48.34 +62 13 35.3 23.2 0′′.42 W, 0′′.1 N
2003er May 25.4 12 36 32.27 +62 07 35.2 23.2 0′′.7 W, 0′′.7 N
2003es May 25.5 12 36 55.39 +62 13 11.9 23.8 0′′.49 W, 0′′.3 N
2003et May 25.7 12 35 55.87 +62 13 32.8 25.3 0′′.5 W, 0′′.14 N
2003eu May 25.7 12 36 05.90 +62 11 01.6 24.7 0′′.7 W, 0′′.3 N
2003ev June 1.3 13 10 31.80 −21 39 49.6 18.5 2′′.2 E, 6′′.4 N

Further magnitudes by the respective observers: SN 2003en, 2002 May 21 UT, [26.0; 2003 Feb. 20.4, 25.9; Apr. 5.5, 26.1; May 21.8, [26.2. SN 2003eo, 2002 May 11.20, [21.0; 2003 Feb. 23.57, [19.9; Apr. 2.14, 20.6; May 6.25, [20.3. SN 2003ep in NGC 7053, 2002 Nov. 30.2, [19.0; 2003 May 28.5, 16.3; June 3.5, 16.7. SN 2003eq, Apr. 6.4, [25.5. SN 2003er, Apr. 3.5, [25.5. SN 2003es, Apr. 3.5, [26.0. SN 2003et, Feb. 22.2, [25.5. SN 2003eu, Feb. 22.2, [26.5. SN 2003ev, Apr. 9.4, [19.0; June 2.3, 18.5.

VARIABLE STAR NEAR MCG +07-36-33 Weisz and Li also report the LOTOSS/KAIT discovery of a possible

supernova (mag ∼ 18.6 on June 1.4 UT, 18.3 on June 2.4, and 18.1 on June 3.4; fainter than 19.0 on May 22.4) located at α = 17^{h}44^{m}07^{s}.72, δ = +40^{o}52′51′′.6 (equinox 2000.0), which is 9′′.5 west and 48′′.9 north of the nucleus of MCG +07-36-33.

2003 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green